# Distributional Vision-Language Alignment by Cauchy-Schwarz Divergence

### v1

@@ Abstract
Multimodal alignment is crucial for various downstream tasks such as cross-modal generation and retrieval. Previous multimodal approaches like CLIP maximize the mutual information mainly by aligning pairwise samples across modalities while overlooking the distributional differences, leading to suboptimal alignment with modality gaps. In this paper, to overcome the limitation, we propose CS-Aligner, a novel and straightforward framework that performs distributional vision-language alignment by integrating Cauchy-Schwarz (CS) divergence with mutual information. In the proposed framework, we find that the CS divergence and mutual information serve complementary roles in multimodal alignment, capturing both the global distribution information of each modality and the pairwise semantic relationships, yielding tighter and more precise alignment. Moreover, CS-Aligher enables incorporating additional information from unpaired data and token-level representations, enhancing flexible and fine-grained alignment in practice. Experiments on text-to-image generation and cross-modality retrieval tasks demonstrate the effectiveness of our method on vision-language alignment.
@@ 1 Introduction
(a) Without CS-Aligner
@@
Figure 1: TSNE visualizations of CLIP text and image features without (a) and with (b) CS-Aligner. The original CLIP feature distributions reveal a clear domain gap (a). Adapting the model with our CS-Aligner effectively eliminates the modality gap, leading to tighter alignment.
Modality alignment is a cornerstone of multimodal representation learning, enabling success across diverse applications such as image-text retrieval (Huang et al., 2024; Koukounas et al., 2024), text-to-image (T2I) generation (Ramesh et al., 2022; Razzhigaev et al., 2023), and multimodal chatbots (Zhu et al., 2023). As a pioneering work in this field, CLIP (Radford et al., 2021) leverages contrastive loss to maximize the mutual information between paired text and image representations, effectively capturing pairwise and semantic relationships. Its versatility has made it a foundation for many multimodal tasks.
Although widely adopted, CLIP suffers from a persistent modality gap between text and image representations in its latent space. As shown in Fig. 1, text and image embeddings often fail to align precisely and may remain separated from each other. This modality gap has been observed and explored in prior studies (Zhou et al., 2023; Liang et al., 2022; Shi et al., 2023), which attribute the issue to factors such as cone effects (Liang et al., 2022) or suboptimal latent space structures (Shi et al., 2023). Intriguingly, Liang et al. (2022) observed a phenomenon that CLIP’s contrastive learning objective may inadvertently exacerbate this gap, implying mutual information alone is insufficient for aligning text and image representation distributions.
@@
Several strategies have been proposed to address the modality gap, such as projection modules with cosine similarity (Zhou et al., 2023; Gao et al., 2024; Huang et al., 2024) and geodesic multimodal mixup (Oh et al., 2024). UnCLIP-based models like DALL-E 2 (Ramesh et al., 2022) and Kandinsky (Razzhigaev et al., 2023) employ text-to-image prior modules (e.g., diffusion models) to map text embeddings to image feature space. A more recent alternative Eclipse (Patel et al., 2024) uses $\ell_{2}$ loss to train a prior adapter. These works aim to transform representations across modalities for alignment. However, they remain exploring alignment sample-wisely and heavily rely on pairwise data. Although sample-wise alignment effectively captures semantic information, it falls short in aligning entire data distributions. Similar to the InfoNCE in CLIP, the methods struggle to match the representation spaces across modalities, ultimately limiting the overall alignment. Moreover, the reliance on carefully curated text-image pairs limits scalability and applicability to real-world scenarios with unpaired and noisy datasets (Lin et al., 2014; Li et al., 2023).
@@
To address these challenges, we propose CS-Aligner, a novel distributional approach that incoporates Cauchy-Schwarz (CS) divergence (Principe et al., 2000b) for modality alignment. As a symmetric measure, CS divergence robustly and efficiently estimates the distance between any representation distributions without parametric distributional assumptions, making it highly suitable for multimodal distribution alignment. Furthermore, we also analyze the complementary roles of CS divergence and mutual information in multimodal alignment and propose integrating these two metrics within CS-Aligner. This enables CS-Aligner to align vision and language representations in both distributional and sample-wise levels, considering both the global modality and local semantic information, leading to more comprehensive and tighter alignment as shown in Fig. 1.
@@
Moreover, based on the distributional approach, CS-Aligner enables alignment with additional unpaired data, such as (a) single images with multiple captions or (b) entirely unpaired vision-language data, introducing more distributional information from richer, unstructured multimodal data for alignment robustness and flexibility in real-world scenarios. Beyond the unpaired alignment, we also introduce a novel token-level alignment scheme for vision-language models, which integrates more detailed information in diverse tokens to enhance multimodal alignment. Extensive experiments on downstream tasks, including T2I generation and image-text retrieval, demonstrate the effectiveness of our approach.
2 Related work
Vision-language alignment and applications. CLIP (Radford et al., 2021) serves as a foundational model for vision-language alignment in multimodal tasks. Several works have enhanced CLIP through techniques such as momentum distillation (Li et al., 2021) and noisy text supervision (Jia et al., 2021). Despite its success, CLIP suffers from a persistent modality gap between text and image representations. Prior studies (Zhou et al., 2023; Liang et al., 2022; Shi et al., 2023) attribute this gap to factors such as cone effects (Liang et al., 2022) and suboptimal latent space structures (Shi et al., 2023). To address this, various strategies have been proposed, including projection adapters (Zhou et al., 2023; Gao et al., 2024; Huang et al., 2024), geodesic multimodal mixup (Oh et al., 2024), and parameter-efficient fine-tuning (Zanella & Ben Ayed, 2024). Recent works also improve CLIP by large language models (LLMs) (Jang et al., 2024; Koukounas et al., 2024; Huang et al., 2024) for downstream tasks such as image-text retrieval.
In addition to image-text retrieval, text-to-image (T2I) generation is another application that reflects the vision-language alignment capability. T2I has advanced significantly over the past decades, driven by both diffusion-based (Ramesh et al., 2021; Rombach et al., 2022; Saharia et al., 2022; Nichol et al., 2021) and GAN-based models (Zhang et al., 2017; Tao et al., 2023). Among diffusion-based methods, the unCLIP framework (Ramesh et al., 2021, 2022) employs a two-stage architecture with a CLIP-guided diffusion prior and a decoder (e.g., DALL-E-2 (Ramesh et al., 2022) or Karlo (Donghoon et al., 2022)). Its prior module $g_{\phi}$ maps text representations $\mathbf{y}$ to image ones $\mathbf{x}$ by a diffusion model. Recently, Eclipse (Patel et al., 2024) employs an $\ell_{2}$ loss to simplify the prior loss by eliminating diffusion time and intruding a noise $\epsilon$ term:
However, these methods still rely on pairwise loss (e.g., $\ell_{2}$). In contrast, our approach introduces distributional alignment for a more holistic modality alignment.
Cauchy-Schwarz divergence. The Cauchy-Schwarz (CS) divergence (Principe et al., 2000a, b) is derived from the Cauchy-Schwarz inequality for square-integrable functions. It serves as a symmetric distribution distance metric with notable properties, such as the ability to measure conditional distributions (Yu et al., 2023) and the closed-form expression for mixtures of Gaussians (Kampa et al., 2011). CS divergence has been successfully applied across various domains, including deep clustering (Trosten et al., 2021), disentangled representation learning (Tran et al., 2022), and deep regression (Yu et al., 2024). Moreover, due to its advantage of estimating discrepancy between conditional distributions, it has demonstrated success in the domain adaption area (Yin et al., 2024) and time series clustering (Yu et al., 2023). However, the utility of CS divergence in foundation models remains unclear and unexplored.
@@ 3 Methodology
(a) High MI & Low Divergence.
(b) High MI & High Divergence.
(c) Low MI & Low Divergence.
Figure 2: Toy examples: mutual information and distribution divergence between two data distributions. Distributions with the same high mutual information value can exhibit either small (a) or large (b) distributional distances, demonstrating that mutual information alone is insufficient for multimodal alignment. In addition, distribution divergence measures the closeness between distributions but does not guarantee that the underlying random variables are statistically correlated (c).
In this section, we first review the mutual information used in previous multimodal methods and analyze its limitations for alignment. Then we introduce the novel CS-Aligner framework for distributional multimodal alignment and detail the estimations of its terms. After that, we extend the method by incorporating additional distribution information. Finally, we provide parameter-efficient implementations.
3.1 Mutual Information is insufficient for alignment
Previous multimodal methods like CLIP (Radford et al., 2021) learn text and image representations in a shared space by maximizing lower bounds (e.g., InfoNCE (Oord et al., 2018)) of mutual information between modalities:
where $p(\mathbf{x})$ and $p(\mathbf{y})$ are the distributions of image and text features. $p(\mathbf{x},\mathbf{y})$ denotes the joint probability. This objective is optimized via the InfoNCE bound (Oord et al., 2018), which approximates $I(\mathbf{x};\mathbf{y})$ using paired data samples $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$. We denote image-text pairs as $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$. The multimodal InfoNCE loss combines symmetric image-text and text-image alignment terms:
$\mathcal{L}_{\text{InfoNCE}}$ $=\frac{1}{2}\left(\mathcal{L}_{\text{I2T}}+\mathcal{L}_{\text{T2I}}\right),$ (3)

### v2

# Distributional Vision Language Alignment by Cauchy-Schwarz Divergence

@@ Abstract
Multimodal alignment is crucial for various downstream tasks such as cross-modal generation and retrieval. Previous multimodal approaches like CLIP utilize InfoNCE to maximize mutual information, primarily aligning pairwise samples across modalities while overlooking distributional differences. In addition, InfoNCE has inherent conflict in terms of alignment and uniformity in multimodality, leading to suboptimal alignment with modality gaps. To overcome the limitations, we propose CS-Aligner, a novel framework that performs distributional vision-language alignment by integrating Cauchy-Schwarz (CS) divergence with mutual information. CS-Aligner captures both the global distribution information of each modality and the pairwise semantic relationships. We find that the CS divergence seamlessly addresses the InfoNCE’s alignment-uniformity conflict and serves complementary roles with InfoNCE, yielding tighter and more precise alignment. Moreover, by introducing distributional alignment, CS-Aligner enables incorporating additional information from unpaired data and token-level representations, enhancing flexible and fine-grained alignment in practice. Experiments on text-to-image generation and cross-modality retrieval tasks demonstrate the effectiveness of our method on vision-language alignment.
@@ 1 Introduction
Modality alignment aims to map the paired modal inputs (e.g., text and image) into a shared feature space, enabling success across diverse applications such as image-text retrieval [1, 2], text-to-image (T2I) generation [3, 4], and multimodal chatbots [5]. As a pioneering work in this field, CLIP [6] leverages InfoNCE loss (a.k.a. contrastive loss) to maximize the mutual information between paired text and image representations, effectively capturing pairwise and semantic relationships. Its versatility has made it a foundation for many multimodal tasks [3, 7].
Despite its success, CLIP and its variants [8, 9] exhibit a persistent modality gap, a misalignment between text and image representations in the shared latent space. As shown in Fig. 1, text and image embeddings often fail to align precisely and may remain separated from each other. This phenomenon has been widely observed [10, 11, 12] and is attributed to issues such as cone effects [11] or suboptimal latent space geometry [12]. Intriguingly, Liang et al. [11] observed that CLIP’s InfoNCE loss could inadvertently exacerbate the modality gap, since, as analyzed in Sec. 2, InfoNCE loss can be decomposed into alignment and uniformity components, which indeed conflict with each other during multimodal alignment.
(a) InfoNCE
@@
Figure 1: TSNE visualizations of CLIP text and image features without (a) and with (b) CS-Aligner. The original CLIP feature distributions reveal a clear domain gap (a). Adapting the model with our CS-Aligner effectively eliminates the modality gap, leading to tighter alignment (b). Consequently, CS-Aligner yields a lower overall $\ell_{2}$ distance between paired image-text features (c).
@@
Several strategies have been proposed to mitigate the modality gap, such as projection modules with cosine similarity [10, 13, 1] and geodesic multimodal mixup [14]. UnCLIP-based models like DALL-E 2 [3] and Kandinsky [4] employ text-to-image prior modules (e.g., diffusion models) to map text embeddings to image feature space for alignment. A more recent alternative Eclipse [15] uses $\ell_{2}$ loss to train a prior adapter for text and image alignment. These works aim to transform representations across modalities for alignment. However, they explore alignment sample-wisely, heavily relying on pairwise data. Although sample-wise alignment effectively captures semantic information, it falls short in aligning entire data distributions. Similar to the InfoNCE in CLIP, the methods struggle to match the representation spaces across modalities, ultimately limiting the overall alignment. The reliance on carefully curated text-image pairs also limits the scalability and applicability to real-world scenarios with unpaired and noisy datasets [16, 17]. Moreover, the theoretical conflict of InfoNCE for multimodal alignment is still under exploration.
@@
To address these challenges, we propose CS-Aligner, a novel distributional approach that incorporates Cauchy-Schwarz (CS) divergence [18] for modality alignment. As a symmetric measure, CS divergence robustly and efficiently estimates the distance between any representation distributions without parametric distributional assumptions, making it highly suitable for multimodal distribution alignment. Furthermore, we analyze the inherent conflict between alignment and uniformity in InfoNCE within the multimodal setting, and find that CS divergence can effectively mitigate the conflict and remains compatible with InfoNCE by estimating via a kernel density estimator (KDE) [19]. This enables CS-Aligner to align vision and language representations in both distributional and sample-wise levels, considering both the global modality and local semantic information, leading to more comprehensive, consistent, and tighter alignment as shown in Fig. 1  and Fig. 1.
@@
Moreover, the distributional nature of CS-Aligner enables alignment with unpaired multimodal data, including cases where a) a single image is associated with multiple captions, or b) vision and language inputs are entirely unpaired. This flexibility allows our method to leverage rich and unstructured datasets and improve alignment robustness beyond curated benchmarks. Beyond unpaired alignment, we introduce a token-level alignment strategy, which further enriches the multimodal representation by aligning fine-grained visual and textual tokens, enhancing the semantic precision of the learned embeddings. Extensive experiments on downstream tasks, including T2I generation and image-text retrieval, demonstrate the effectiveness of our approach.
2 InfoNCE is insufficient for alignment
Previous multimodal methods like CLIP [6] learn text and image representations in a shared space by maximizing lower bounds (e.g., InfoNCE [20]) of mutual information between modalities:
where $p(\mathbf{x})$ and $p(\mathbf{y})$ are respectively the distributions of image and text features, and $p(\mathbf{x},\mathbf{y})$ denotes their joint probability. Although widely used, it suffers from two limitations.
Figure 2: Toy examples: mutual information (MI $\uparrow$) and distribution divergence ($\downarrow$) between two distributions. Distributions with the same high mutual information value can exhibit either large (a) or small (b) distributional distances, demonstrating that MI alone is insufficient for multimodal alignment. Moreover, distribution divergence measures the closeness between distributions but does not guarantee that the underlying random variables are statistically correlated (c).
Limitation1: Mutual information is insufficient for multimodal alignment. Although widely adopted, mutual information alone is insufficient for effective modality alignment [11]. The reason is that mutual information quantifies the statistical dependence between two random variables [21], ensuring correlation maximization between two random variables. However, it does not guarantee that the distributions $p(\mathbf{x})$ and $p(\mathbf{y})$ are statistically similar or close to each other in terms of their underlying distributions. In other words, the embedding distributions of two modalities can differ significantly or be far apart, yet exhibit strong dependence.
We illustrate this issue using a toy example in Fig. 2. Fig. 2a shows that despite strong dependence and high mutual information, the representation distributions of two representations or random variables can remain misaligned and be far from each other, resulting in a high divergence. This issue is also observed in the CLIP model pretrained with InfoNCE, where the vision and language representations exhibit a noticeable distributional gap, as shown in Fig. 1. This gap results in inconsistently aligned multimodal features, hindering the clear representation of shared semantics and disrupting effective mapping between modalities. Ultimately, this misalignment degrades performance in downstream tasks, including cross-modality generation. Ideally, the desired multimodal representations should be highly correlated with low distributional divergence, as depicted in Fig. 2b. Notably, although directly minimizing the divergence between distributions may reduce the distributional gap, it risks creating independent multimodal distributions without common semantic information (Fig. 2c). Therefore, maximizing mutual information and minimizing divergence complement each other to achieve effective multimodal representation alignment. Details are provided in Appendix B.
Limitation2: InfoNCE includes conflicting terms for multimodal alignment. In practice, mutual information is often optimized via minimizing the InfoNCE loss [20] which estimates $I(\mathbf{x};\mathbf{y})$ using paired image-text data $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ and contains image-text and text-image alignment terms:
$\mathcal{L}_{\text{InfoNCE}}$ $=-\frac{1}{2N}\sum_{i=1}^{N}\left(h(\mathbf{x}_{i},\mathbf{y}_{i})+h(\mathbf{y}_{i},\mathbf{x}_{i})\right),\quad\ \ h(\mathbf{x},\mathbf{y})=\log\frac{\exp\left(\text{sim}(\mathbf{x},\mathbf{y})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{x},\mathbf{y}_{j})/\tau\right)},$ (2)
where $\text{sim}(\cdot,\cdot)$ is cosine similarity and $\tau$ is temperature. Critically, the InfoNCE loss in Eq. (2) requires paired data $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i}^{N}$, and cannot work under unpaired setting.
As analyzed in [22], the InfoNCE loss can be decomposed as the sum of the alignment and uniformity terms i.e., $\mathcal{L}_{\text{InfoNCE}}\approx\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{uniform}}$:
[…]
where $t$ and $\alpha$ are hyperparameters. $p_{\text{pair}}$ denotes the image-text pairs distribution. Minimizing $\mathcal{L}_{\text{align}}$ encourages pairwise alignment. In unimodality, minimizing $\mathcal{L}_{\text{uniform}}$ promotes representations that are uniformly distributed on the unit hypersphere, a desirable property for representation learning [22]. However, in multimodal alignment, $\mathcal{L}_{\text{uniform}}$ may conflict with $\mathcal{L}_{\text{align}}$.
Remark 2.1.
The uniformity and alignment terms in InfoNCE conflict with each other in multimodal alignment. Applying Taylor expansions ($\mathbb{E}(e^{-\mathbf{x}})\approx 1-\mathbb{E}(\mathbf{x})$ and $\log(1-\mathbf{x})\approx-\mathbf{x}$) on $\mathcal{L}_{\text{uniform}}$, the uniformity term becomes:
$\mathcal{L}_{\text{uniform}}\approx-t\mathbb{E}_{\mathbf{x},\mathbf{y}\sim p(\mathbf{x},\mathbf{y})}\left[||\mathbf{x}-\mathbf{y}||_{2}^{2}\right]=-t\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}+p_{\text{unpair}}}\left[||\mathbf{x}-\mathbf{y}||_{2}^{2}\right],$ (4)
where $p(\mathbf{x},\mathbf{y})=p_{\text{pair}}+p_{\text{unpair}}$, and $p_{\text{unpair}}$ denotes the distribution of unpaired image and text. Consequently, the combination of the two (InfoNCE) can be written as:
$\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{uniform}}\approx\mathbb{E}_{(x,y)\sim p_{\text{pair}}}\left[||\mathbf{x}-\mathbf{y}||_{2}^{\alpha}\right]-t\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}+p_{\text{unpair}}}\left[||\mathbf{x}-\mathbf{y}||_{2}^{2}\right].$ (5)
The alignment contribution ($\mathcal{L}_{\text{align}}$) in Eq. (3) can be largely suppressed or even canceled (when $t=1$) due to the opposing term in Eq. (5), leaving only negative pairs influential. Essentially, $\mathcal{L}_{\text{align}}$ promotes alignment across modalities, whereas $\mathcal{L}_{\text{uniform}}$ encourages dissimilarity among negative pairs without preserving intra-modal structure. This inherent conflict can result in local minima, driving alignment and uniformity in opposing directions and ultimately leading to a modality gap. Thus, InfoNCE alone may lead to suboptimal alignment between modalities.
@@ 3 Methodology
In this section, we address the incapability of mutual information on aligning distributions and the conflicts in InfoNCE for multimodal alignment. To this end, we first introduce a novel distributional multimodal alignment framework, CS-Aligner. Then, we analyze that with the KDE, the proposed method is able to address the uniformity-alignment conflicts of InfoNCE. Finally, we extend CS-Aligner to the unpaired data, including token-level alignment.
3.1 CS-Aligner: distributional multimodal alignment
To mitigate limitation 1 in Sec. 2, we explicitly minimize the distribution divergence between $p(\mathbf{x})$ and $p(\mathbf{y})$. In practice, $p(\mathbf{x})$ and $p(\mathbf{y})$ may follow arbitrary distributions with minimal intersection, which may often occur in the multimodal setting. Hence, a robust divergence metric must accommodate unpredictable variability and limited support overlap for effective distribution alignment.
To this end, we propose a distributional alignment framework, namely CS-Aligner, which leverages the Cauchy-Schwarz divergence ($D_{\text{CS}}$[23, 18], as illustrated in Fig. 3. The objective is:
$\min-I(\mathbf{x};\mathbf{y})+\lambda D_{\text{CS}}(p(\mathbf{x}),p(\mathbf{y})),$ (6)
where $\lambda$ is a hyperparameter balancing the mutual information term and the divergence penalty. CS divergence, $D_{\text{CS}}$, is a symmetric and robust metric to quantify the distance between any two probability density functions $p$ and $q$, defined over the same support $\omega$ as:
[…]
The CS divergence satisfies $0\leq D_{\text{CS}}<\infty$, and equals zero if and only if $p=q$. By introducing $D_{\text{CS}}$ in Eq. (6), instead of solely minimizing pairwise distance, our method also aligns the distributions of modalities, leading to more robust and efficient multimodal alignment, as shown in Fig. 3.
CS divergence estimation. To estimate CS divergence, we introduce non-parametrical KDE [19], eliminating the need for explicit parametric assumptions about the underlying distributions. This provides significant flexibility in measuring distributional distance. Given i.i.d. samples $\{\mathbf{x}_{i}\}_{i=1}^{M}\sim p(\mathbf{x})$ and $\{\mathbf{y}_{i}\}_{i=1}^{N}\sim p(\mathbf{y})$, the empirical CS divergence estimator is given by [24]:
$\widehat{D}_{\text{CS}}(p(\mathbf{x});p(\mathbf{y}))\!=\!\log\!\Big{(}\!\frac{1}{M^{2}}\!\!\sum_{i,j=1}^{M}\!\!\!\kappa({\mathbf{x}}_{i},{\mathbf{x}}_{j})\Big{)}\!+\log\!\Big{(}\!\frac{1}{N^{2}}\!\!\sum_{i,j=1}^{N}\!\!\!\kappa({\bf y}_{i},{\mathbf{y}}_{j})\Big{)}-2\log\!\Big{(}\!\frac{1}{MN}\!\sum_{i=1}^{M}\!\sum_{j=1}^{N}\!\!\kappa({\mathbf{x}}_{i},{\mathbf{y}}_{j})\Big{)}.$ (8)
where $\kappa$ is a kernel function such as Gaussian $\kappa_{\sigma}(\mathbf{x},\mathbf{y})=\exp(-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/2\sigma^{2})$ with kernel width $\sigma$. This estimator is symmetric, differentiable, and computationally efficient, making it suitable for multimodal alignment. Moreover, the third term in Eq. (8) ensures that $\widehat{D}_{\text{CS}}(p(\mathbf{x});p(\mathbf{y}))\to\infty$ only when $\mathbb{E}(\kappa({\mathbf{x}},{\mathbf{y}}))\to 0$ (i.e., when the distributions do not overlap). However, as long as there is a nonzero overlap between the distributions, the estimator remains well-defined and valid.
Hence, CS-Aligner remains reliable even when the two distributions initially have limited overlap, a common scenario in multimodal tasks. Additionally, its symmetry and non-parametric estimation properties ensure consistent and unbiased multimodal alignment. Consequently, our method ensures both semantic and distributional alignment, enabling robust and efficient multimodal learning.
When estimating the mutual information term $I(\mathbf{x},\mathbf{y})$ via InfoNCE (Eq. (2)), unlike other distribution divergences, CS divergence effectively addresses InfoNCE’s inherent alignment-uniformity conflict.
Uniformity and Alignment with CS Divergence. Using the Gaussian kernel $\kappa_{t}(\mathbf{x},\mathbf{y})=\exp(-t\|\mathbf{x}-\mathbf{y}\|_{2}^{2})$ for CS divergence and combining the alignment and uniformity components of InfoNCE, the full objective of Eq. (6) can be expressed as
$\mathcal{L}$ $=\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\mathrm{pos}}}\bigl{[}\|\mathbf{x}-\mathbf{y}\|_{2}^{\alpha}\bigr{]}+\log\,\mathbb{E}_{\bf{x}\sim p(\mathbf{\bf{x}}),\,\bf{y}\sim p(\mathbf{\bf{y}})}\bigl{[}\kappa_{t}(\mathbf{x},\mathbf{y})\bigr{]}$ (9)
[…]
When $\lambda=1$, this reduces to the following alignment–uniformity decomposition: